# FlowBot++: Learning Generalized Articulated Objects Manipulation via Articulation Projection

### v1

@@ 4.1 A New Representation of Articulated Objects
In order to overcome the limitations of previous work, if we could estimate the articulation parameters $\vec{v}$ (or equivalently, $\boldsymbol{\omega}$ and $v$), then we would be able to derive the full trajectory of how any arbitrary point $p$ on the object would move under articulation of the object under kinematic constraints of $\vec{v}$. To achieve this, we formally define a 3D dense visual representation of articulation parameters on top of Articulation Flow; we illustrate each representation graphically in Fig. 2. In this work, we introduce another 3D representation that densely represents the object’s articulation parameters, Articulation Projection - for each point on the articulated part of the object, we define a vector $r_{p}$ that represents the displacement from the point itself to its projection onto the part’s articulation axis $\vec{v}(t)$. For each point $p\in\mathbb{R}^{3}$ of articulated object $\mathcal{P}$, we define the Articulation Projection $r_{p}$ mathematically as follows:
@@ 4.3 Jointly Learning Articulation Flow and Articulation Projection
Empirically, the Articulation Projection prediction is sometimes less accurate than the Articulation Flow prediction. On the other hand, the Articulation Flow represents the instantaneous motion and thus needs to predicted at each timestep, which prevents multi-step planning. To obtain the both of both worlds, we correct the Articulation Projection $r_{p}$ prediction with the flow estimate using Gram-Schmidt on $r_{p}$ to make it perpendicular to the Articulation Flow $f_{p}$; the correction is computed as: $\tilde{r}_{p}=r_{p}-\mathrm{proj}_{f_{p}}r_{p}$, which is used in place of $r_{p}$ in Eq. 3. The effect of using Gram-Schmidt is shown in Fig. 4a.
@@ 5 Experiments
UMP-DI [12] 0.29 0.32 0.33 0.16 0.18 0.37 0.14 0.19 0.28 0.72 0.00 0.55 0.36 0.32 0.62 0.15 0.00 0.41 0.61 0.17 0.34 0.38 0.58
Normal Direction 0.40 0.52 0.67 0.16 0.19 0.51 0.60 0.13 0.11 0.55 0.61 0.32 0.39 0.69 0.57 0.04 0.00 0.67 0.19 0.66 1.00 0.43 0.26
Screw Parameters [1] 0.40 0.42 0.40 0.42 0.18 0.57 0.45 0.27 0.59 0.51 0.58 0.06 0.18 0.19 0.26 0.08 0.08 0.23 0.06 0.14 0.21 0.34 0.24
DAgger E2E [31] 0.64 0.39 0.85 0.61 0.73 0.50 1.00 0.96 0.90 0.54 0.48 0.10 0.83 0.73 0.62 1.00 0.80 0.95 0.73 0.83 0.98 0.81 0.85
FlowBot3D (AF Only) [6] 0.12 0.32 0.23 0.11 0.09 0.02 0.00 0.09 0.32 0.04 0.13 0.00 0.15 0.00 0.21 0.00 0.00 0.33 0.22 0.09 0.07 0.19 0.35
@@ Acknowledgments
If a paper is accepted, the final camera-ready version will (and probably should) include acknowledgments. All acknowledgments go at the end of the paper, including thanks to reviewers who gave useful comments, to colleagues who contributed to the ideas, and to funding agencies and corporate sponsors that provided financial support.
@@ Appendix C Baselines
  • UMP-DI: We implement UMPNet’s Direction Inference network (DistNet) [12], where instead of bootstrapping an action scoring function from interaction, we learn the scoring function by regressing the cosine distance between a query vector and the ideal flow vector for a contact point. At test time, we select the contact point based on ground-truth 3DAF, and after contact has been achieved we use CEM to optimize the scoring function to predict the action direction at every timestep.
@@
UMP-DI [12] 0.52 0.60 0.33 0.65 0.73 0.29 0.67 0.80 0.50 0.11 1.00 0.00 0.45 0.83 0.03 0.50 1.00 0.31 0.29 0.78 0.33 0.31 0.20
Normal Direction 0.31 0.40 0.00 0.51 0.71 0.00 0.00 0.80 0.50 0.00 0.00 0.50 0.31 0.00 0.00 0.50 1.00 0.00 0.55 0.00 0.00 0.10 0.64
Screw Parameters [1] 0.50 0.50 0.53 0.51 0.80 0.21 0.55 0.60 0.17 0.37 0.43 0.80 0.67 0.17 0.63 0.67 0.92 0.69 0.92 0.83 0.75 0.50 0.72
DAgger E2E [31] 0.14 0.60 0.00 0.26 0.09 0.28 0.00 0.00 0.00 0.00 0.25 0.00 0.04 0.00 0.00 0.00 0.20 0.00 0.17 0.02 0.00 0.00 0.00
FlowBot3D (AF Only) [6] 0.77 0.57 0.56 0.88 0.82 0.86 1.00 0.80 0.50 0.78 0.75 1.00 0.69 1.00 0.56 1.00 1.00 0.38 0.43 0.84 0.83 0.43 0.44
@@
FlowBot++ (Ours - Combined) 0.79 0.61 0.59 0.89 0.82 0.96 0.86 0.78 0.52 0.78 0.75 1.00 0.74 1.00 0.68 1.00 1.00 0.43 0.63 0.88 0.81 0.45 0.52

### v3

@@ 4.1 A New Representation of Articulated Objects
In order to overcome the limitations of previous work [6, 12], if we could estimate the articulation parameters $\vec{v}$ (or equivalently, $\boldsymbol{\omega}$ and $v$), then we would be able to derive the full trajectory of any arbitrary point $p$ on the object would move under articulation of the object under kinematic constraints of $\vec{v}$. To achieve this, we formally define a 3D dense visual representation of articulation parameters on top of Articulation Flow; we illustrate each representation graphically in Fig. 2. In this work, we introduce another 3D representation that densely represents the object’s articulation parameters, Articulation Projection - for each point on the articulated part of the object, we define a vector $r_{p}$ that represents the displacement from the point itself to its projection onto the part’s articulation axis $\vec{v}(t)$. For each point $p\in\mathbb{R}^{3}$ of articulated object $\mathcal{P}$, we define the Articulation Projection $r_{p}$ mathematically as follows:
@@ 4.3 Jointly Learning Articulation Flow and Articulation Projection
Empirically, the Articulation Projection prediction is sometimes less accurate than the Articulation Flow prediction. On the other hand, the Articulation Flow represents the instantaneous motion and thus needs to predicted at each timestep, which prevents multi-step planning. To obtain the best of both worlds, we correct the Articulation Projection $r_{p}$ prediction with the flow estimate using Gram-Schmidt on $r_{p}$ to make it perpendicular to the Articulation Flow $f_{p}$; the correction is computed as: $\tilde{r}_{p}=r_{p}-\mathrm{proj}_{f_{p}}r_{p}$, which is used in place of $r_{p}$ in Eq. 3. The effect of using Gram-Schmidt is shown in Fig. 4a.
@@ 5 Experiments
UMP-Net [12] 0.18 0.18 0.17 0.32 0.32 0.05 0.06 0.12 0.24 0.23 0.18 0.08 0.15 0.23 0.14 0.04 0.00 0.25 0.27 0.09 0.21 0.13 0.19
Normal Direction 0.41 0.52 0.67 0.16 0.19 0.51 0.60 0.13 0.14 0.45 0.71 0.31 0.40 0.79 0.43 0.07 0.00 0.64 0.15 0.70 1.00 0.41 0.29
Screw Parameters [1] 0.42 0.44 0.42 0.44 0.16 0.64 0.52 0.21 0.57 0.43 0.58 0.11 0.24 0.28 0.25 0.14 0.12 0.26 0.12 0.11 0.12 0.32 0.26
DAgger E2E [31] 0.52 0.25 0.86 0.51 0.71 0.55 1.00 0.86 0.80 0.74 0.58 0.40 0.64 0.53 0.52 1.00 0.65 0.75 0.54 0.63 0.68 0.87 0.74
FlowBot3D (AF Only) [6] 0.17 0.42 0.22 0.16 0.17 0.03 0.00 0.20 0.51 0.07 0.00 0.08 0.21 0.17 0.29 0.00 0.06 0.21 0.10 0.06 0.16 0.29 0.73
@@ Acknowledgments
This material is based upon work supported by the National Science Foundation under Grant No. IIS-2046491. We are grateful to Prof. Daniel Seita for his helpful feedback and discussion on the paper.
@@ Appendix C Baselines
  • UMP-Net: This is the same implementation/code base as provided in [12].
@@
UMP-Net [12] 0.73 0.73 0.71 0.60 0.49 0.89 0.90 0.79 0.60 0.64 0.78 0.86 0.75 0.55 0.80 0.89 1.00 0.66 0.64 0.77 0.64 0.75 0.76
Normal Direction 0.51 0.60 0.12 0.62 0.75 0.10 0.10 0.85 0.42 0.12 0.10 0.60 0.45 0.23 0.09 0.60 1.00 0.30 0.51 0.00 0.20 0.22 0.71
Screw Parameters [1] 0.55 0.55 0.58 0.52 0.90 0.20 0.54 0.69 0.19 0.39 0.41 0.85 0.69 0.19 0.73 0.69 0.91 0.63 0.82 0.89 0.85 0.60 0.78
DAgger E2E [31] 0.32 0.63 0.10 0.36 0.19 0.39 0.12 0.13 0.00 0.12 0.38 0.10 0.09 0.00 0.00 0.10 0.25 0.20 0.17 0.02 0.00 0.00 0.00
FlowBot3D (AF Only) [6] 0.81 0.53 0.74 0.81 0.82 0.96 0.99 0.79 0.44 0.90 1.00 0.89 0.70 0.69 0.63 1.00 0.94 0.67 0.89 0.75 0.66 0.69 0.14
@@
FlowBot++ (Ours - Combined) 0.82 0.70 0.62 0.89 0.91 0.96 0.96 0.83 0.72 0.78 0.85 1.00 0.76 1.00 0.68 1.00 1.00 0.43 0.63 0.91 0.81 0.45 0.82